For an airplane in flight, the average speed of the air relative to the plane is 317 m/s below the wings and 348 m/s above the wings. The bottom of each wing has a surface area of 29.9 m2. Calculate the lift, which is the net upward force of air on the plane.
Answer #1

vb =317 m/s , va =348 m/s

Ab = 29.9 m^2

density of air d = 1.2 kg/m^3

Volume flow rate is constant

vb*Ab = va*Aa

Aa = vb*Ab/va = (317*29.9)/348

Aa = 27.24 m^2

F =PA

P = (1/2)dv^2

Net force Fb -Fa = (1/2)dvb2Ab - (1/2)dva2Aa

= (1/2)(1.2) [ (348*348*27.24) - (317*317*29.9)]

Fnet =1.77x105 N